I need to solve this by the addition method.
3/7x+5/9y=27
1/9x+2/7y=7

Answer by Alan3354(69443)

I need to solve this by the addition method.
3/7x+5/9y=27
1/9x+2/7y=7
------------------
Assuming you mean
(3x/7) + (5y/9) = 27
(x/9) + (2y/7) = 7
Multiply thru both by 63 to eliminate fractions.
27x + 35y = 1701
7x + 18y = 441
Some big numbers, but easier to deal with than fractions.
To get the same coefficients for x, multiply the 1st eqn by 7, the 2nd by 27
189x + 245y = 11907
189x + 486y = 11907
-------------- Uhoh
Subtract 2 from 1
0x - 241y = 0
y = 0
x = 63
It does check, and fits both eqns